# DreamDrone

### v2

@@ 5 Conclusion
In this work, we pioneer the concept of zero-shot, training-free text-to-perpetual view generation, introducing DreamDrone for generating geometry-consistent scenes. We harness the robust capabilities of a pre-trained text-to-image diffusion model and develop a novel feature-correspondence-guidance diffusion process for creating novel views. Our quantitative and qualitative analyses underscore DreamDrone’s superiority over other state-of-the-art methods. We anticipate that our approach will inspire further innovation in AI-Generated Content (AIGC) and make advanced research accessible to a broader spectrum of researchers.
Supplementary Material
@@ 6 Implementation details
We take Stable Diffusion [34] with the pre-trained weights from version 2.1https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the basic text-to-image diffusion and MiDas [32] with weights $\mathrm{dpt\_beit\_large\_512}$https://github.com/isl-org/MiDaS. The overall diffusion timesteps is $1000$. We wrap the latent code at timestep $t_{1}$=$21$ and add more degrees of noise to timestep $t_{2}$=$441$. The threshold $\sigma$ for low-pass wrapping is $20$ and the hyper-parameter $\lambda$ for feature-correspondence guidance is $300$. We conducted the experiments on Titan-RTX GPU. The generated speed is roughly 30 seconds per frame.
@@ 7 Ablation studies
In this section, we conduct our ablation studies on each proposed module. First, we explore the impact of wrapping the latent code at timestep $t_{1}$ in our pipeline on the generation results, with comparative results shown in Fig. 6. According to the principles of diffusion, a smaller timestep means that the corresponding latent code has less noise, thus retaining more complete content and detailed information. When we wrap the latent code at a smaller timestep (for example,$t_{1}=1$ or $21$), it is observed that, with the camera’s movement, DreamDrone still maintains its complete content information quite well. As the timestep increases, with more noise, the content information gets somewhat disrupted, and wrapping the latent code then does not yield as favorable results. This experiment demonstrates that the latent code corresponding to a smaller timestep possesses richer geometry and content information.
Figure 6: Ablation study on the wrapping timestep $t_{1}$.
Next, we investigate how editing the latent code in different ways can enhance the generation results. We set various thresholds for the low-pass wrapping module, where a smaller threshold indicates the retention of less low-frequency information and more high-frequency information. The comparative results are detailed in Fig. 7. The first and second rows display the generation results corresponding to smaller thresholds. It can be observed that setting an excessively small threshold retains too much ineffective or noisy high-frequency information, leading to generated images with an unrealistic abundance of high-frequency details. Conversely, setting the threshold too high, or removing the low-pass wrapping module, results in a lack of retained detail and very blurry images.
Figure 7: Ablation study on the threshold of the low-pass wrapping module $\sigma$.
After obtaining the wrapped latent code, we next investigate the impact of setting $t_{2}$ on the generation results. The size of $t_{2}$ influences the freedom in the denoising process; a larger $t_{2}$ means adding more noise to the wrapped latent code, thereby gaining higher degrees of freedom. The generation results corresponding to different $t_{2}$ settings are shown in Fig. 8. A smaller $t_{2}$ (for example, $t_{2}=161$ or 241) results in insufficient noise added to the wrapped latent code, leading to increasingly blurred generated images. Setting a larger $t_{2}$ produces clearer images with richer details, but an excessively high degree of freedom can decrease the geometric consistency between adjacent frames, adversely affecting their coherence.
Figure 8: Ablation study on the denoising timestep $t_{2}$.
To ensure geometric consistency between adjacent frames during the denoising process, we introduced a feature-correspondence guidance strategy. Different guidance weights lead to significant variations in generation results, as detailed in Fig. 9. When feature-correspondence guidance is not introduced (i.e., $\lambda=0$), the generated image sequence maintains content consistency, but some details may not be consistent. The introduction of feature-correspondence guidance ensures the consistency of details between adjacent frames; however, excessively high guidance weights can impact the realism of the generated images.
Figure 9: Ablation study on the weight parameter of feature-correspondence guidance $\lambda$.
@@ 8 Additional qualitative comparisons
Figure 10: Qualitative comparisons of VideoFusion [23] and our DreamDrone.
@@ 9 More visualization results
In this section, we provide more visualization results. Please refer to Figs. 11 and 12 for details. Moreover, please watch our attached videos for a comprehensive evaluation of our proposed DreamDrone.
Figure 11: Visualization results of our DreamDrone.
Figure 12: Visualization results of our DreamDrone.

### v3

@@ 5 Conclusion
In this work, we propose DreamDrone, a novel approach for generating flythrough scenes from textual prompts without the need for training or fine-tuning. Our method explicitly warps the intermediate latent code of a pre-trained text-to-image diffusion model, enhancing the quality of the generated images and the generalization ability. We propose a feature-correspondence-guidance diffusion process and a high-pass filtering strategy to ensure geometric and high-frequency detail consistency. Experimental results indicate that DreamDrone surpasses current methods in terms of visual quality and authenticity of the generated scenes.
Acknowledgement
This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award Number: MOE-T2EP20122-0006).
DreamDrone: Text-to-Image Diffusion Models
are Zero-shot Perpetual View Generators
Supplement Material
[…]
In this supplementary material, we provide more comprehensive ablation studies, comparisons of visual results with text-to-video methods, and additional visual results of our method. Finally, we discuss the limitations and social impact of this approach.
@@ 6 Implementation details
[…]
@@ 7 Ablation studies
Figure 8: Ablation results for the key components. We perform ablation studies by disabling the key components of our method. We illustrate every five frames for each ablation experiment. Please zoom in for better comparisons.
In this section, we first provide additional ablation results for Fig. 8 as referenced from Fig. 3 in the main text. To enhance the robustness of our ablations, we include one example for each experimental setup. The comprehensive results of the ablation study for each component are illustrated in Fig. 8. For quantitative results, please refer to Tab. 1 in the main text. The simplest method for tasks involving infinite scene generation is frame-by-frame image warping. However, this approach is impractical, as is the direct warping of the latent code. Warping images results in non-integer pixel coordinates, which leads to interpolation-induced blurring and distortion. Furthermore, these errors accumulate with each generated frame, causing a significant degradation in quality, with the images becoming progressively blurred, as shown in warp image and warp latent in Fig. 8.
To address these challenges, we introduce DDPM to increase the degrees of freedom for the diffusion model, facilitating the generation of high-quality images (warp latent + DDPM). With the incorporation of DDPM, the CLIP score improves from $0.125$ to $0.308$ for a series of 32 images. Nevertheless, the introduction of DDPM inadvertently affects the semantic consistency between adjacent frames.
To maintain the integrity of image content, and inspired by previous methods [44, 14], we integrate cross-view attention modules into our framework. As demonstrated in warp latent + DDPM + cross attn. in Fig. 8, the consistency of geometries across views is significantly improved compared to warp latent + DDPM. To ensure high-quality image generation while also maintaining consistency across adjacent views, we propose a feature-correspondence guidance strategy. Comparing warp latent + DDPM + guidance with warp latent + DDPM + guidance + cross view attn., it is evident that semantic consistency between adjacent frames is significantly enhanced after incorporating guidance, as indicated by the noticeable improvements in both PSNR and SSIM scores in Table 1 of the main text. To further improve the cross-view consistency of high-frequency details, we have employed high-pass filtering. This approach aids in preserving the high-frequency details of the current frame, thus enhancing the semantic consistency of high-frequency details between consecutive frames. For instance, the enhanced cross-view consistency, particularly of the house on the left side in Fig. 8, illustrates the effectiveness of adding the proposed modules.
Then, we conduct more detailed ablation studies on each proposed module in a Q&A manner.
Q1:
Does DDIM inversion limit the reconstruction fidelity?
Figure 9: DDIM inversion and image reconstruction. We illustrate the pipeline for evaluating the reconstruction performance using DDIM inversion. The left side is the pipeline and the right side is the reconstructed results at different timesteps $t_{1}$. We visualize two different result samples.
Before warping latent, the most important thing is to ensure we can reconstruct the original image without any editing of the intermediate latent code. To this end, we establish a simple experiment. We obtain the intermediate latent code $x_{t_{1}}$ at different timestep $t_{1}$, denoising the noise, and decode the reconstructed image. In  Fig. 9, the left side is the pipeline of this experiment and the right side is the reconstructed images at different timestep $t_{1}$. We take two sample images as examples. From the results, we can figure out that the images can be reconstructed at different timestep $t_{1}$.
Based on the former discussion, regarding the top branch in Fig. 2 in the main text, the image at the current view can be reconstructed. The reconstruction of the top branch is the foundation of the feature-correspondence guidance and cross-view attention.
Q2:
Why do the image sequences become blurrier when generating more images, no matter whether warping the image or warping the latent code?
The $1^{st}$ and $2^{nd}$ rows in Fig. 3 in the main text show that the images become blurrier when generating more images. Besides, the reconstruction results in  Fig. 9 show that DDIM inversion can reconstruct original images if there is no editing for the intermediate latent code, i.e., warping. It is straightforward that when we fly through, in other words, zoom in, the images, the images will become much blurrier. This is because the warping operation leads to non-integer pixel coordinates. Previous SOTA methods [20, 18, 3] train a refiner to add the details and inpainting or outpainting the missing region when the camera is moving. In this paper, we serve the pre-trained text-to-image diffusion model as a ‘refiner’ due to its powerful generation capacity.
Q3:
Why DDPM forward is needed?
Figure 10: Ablation studies for DDPM forward without high-pass filtering. To evaluate the necessity of DDPM forward operation, we conduct ablation experiments based on the simple pipeline. The corresponding ablation results are shown in Fig. 11
Now we analyze the necessity of the DDPM forward process. As illustrated in Fig. 2 in the main text, we further apply DDPM forward after warping the latent code. Comparing the $5^{th}$ and $6^{th}$ rows in Fig. 8, we can figure out that the image quality improves a lot after DDPM is applied. The details are enhanced and there is no distortion. The side effect of DDPM forward is that the correlation between adjacent views degrades because more degrees of freedom are introduced by DDPM.
Figure 11: The visualization results for evaluating DDPM forward without high-pass filtering. We illustrate the results based on the pipeline shown in Fig. 10. We fix the warping timestep $t_{1}$ and generate image sequences with different DDPM forward timesteps $t_{2}$. To facilitate the comparisons, we further illustrate the generation results using our overall pipeline at last. We visualize every five frames per sample.
To evaluate how DDPM affects the generation results, we fix the warping timestep $t_{1}$ and illustrate the generated image sequences with DDPM at various timesteps $t_{2}$. The pipeline of this ablation experiment is shown in Fig. 10. The results are shown in Fig. 11. A smaller $t_{2}$ means less degree of freedom for the diffusion model, which can result in blurring and distortion. As shown in the first two rows in Fig. 11, the generated images become blurrier when generating more images. A proper $t_{2}$ makes the geometry between adjacent views more consistent. In the $3^{rd}$ and $4^{rd}$ rows in Fig. 11, the geometry layout in the image sequences becomes consistent. For instance, the geometry of the house on the left side of the image looks roughly consistent. Moreover, the image quality is satisfied. As $t_{2}$ becomes larger, more random noise are added to the warped latent code. Though the image quality is promising, the consistency degrades a lot. As shown in the $5^{th}$ and $6^{th}$ rows, the consistency across adjacent views is much worse than $3^{th}$ and $4^{th}$ rows.
Figure 12: Ablation studies for high-pass filtering without DDPM forward. To analyze if the DDPM can be replaced by the high-pass filter, we conduct a simple ablation based on this pipeline. The corresponding ablation results are shown in Fig. 13.
This ablation demonstrates that the DDPM forward module with a proper timestep $t_{2}$ improves the image quality. But the consistency is still not satisfied. That’s the reason why we further propose the feature-correspondence guidance strategy.
Moreover, since the high-pass filter preserves details from the previous view, is it possible to remove DDPM forward and only use the high-pass filter? To this end, we further conduct an ablation experiment. The pipeline for this ablation is shown in Fig. 12. In this pipeline, we remove DDPM forward operation and add the high-pass filter when warping the latent code. The experimental results are shown in Fig. 13. We show two results for each $\sigma$. No matter how large the threshold $\sigma$ is, the high-pass filter cannot help to preserve the details from previous view. The reason is that the high-pass filter can only preserve high-frequency details from the previous view, rather than the low-frequency content. Combined with the results shown in Fig. 14, the low-frequency information dominants the content when combining frequencies from different images. As discussed before, the content would become much blurrier when warping the latent code, which motivates us to propose the feature-correspondence guidance strategy.
Figure 13: The visualization results for high-pass filter without DDPM forward. To evaluate if we can preserve high quality from the previous view using the high-pass filter, we remove the DDPM forward module and visualize the image sequences with different threshold $\sigma$. To facilitate the comparisons, we further illustrate the generation results using our overall pipeline at last. We visualize every five frames per sample.
Q4:
Will the combination of low and high frequencies from different images break up the correlation of the frequency of the original image and introduce more errors?
Figure 14: Toy experiments of low and high-frequency combination from different images. Our toy experiments are illustrated on the left side. We combine the low frequency of Elon Musk and the high frequency of Vincent van Gogh’s self-portrait with various threshold $\sigma$. The higher $\sigma$, the more low-frequency of Elon Musk is used. The results with various $\sigma$ are illustrated on the right side. Please zoom in for comparisons.
To evaluate the feasibility of the frequency combination, we conduct a toy experiment, which is shown in Fig. 14. In this experiment, we first obtain the frequency from two different images and combine the frequencies given different threshold $\sigma$. As shown on the right side in Fig. 14, the content of Elon Musk does not change much with different $\sigma$, which demonstrates the feasibility of frequency combination. An extremely small $\sigma$, for instance, $\sigma=10$, introduces excessive details from van Gogh’s portrait.
In this toy experiment, the content of the two images is extremely different. However, regarding the perpetual view generation task, the content of the adjacent view would not be so different. Now we analyze how different $\sigma$ affects the generation results. We apply all the proposed modules in this experiment and change the $\sigma$ value. The comparison results are shown in Fig. 15. As shown in Fig. 15, a small $\sigma$ neglects more low-frequency content from the previous view, which results in an inconsistency between images. A large $\sigma$ introduces less high-frequency details from the previous view. Though looks consistent, the generated images look not as realistic as $\sigma=20$.
Figure 15: Ablation studies on high-pass filter. We apply all the proposed modules with various thresholds $\sigma$ of the high-pass filter.
@@ 8 Additional qualitative comparisons
Figure 16: Qualitative comparisons of VideoFusion [24] and our DreamDrone. We show the visualization results given three prompts and illustrate every five frames for each sample.
@@ 9 More visualization results
In this section, we provide more visualization results. We generate 120 images for each prompt and visualize one image from every third frame. Please refer to Figs. 17 and 18 for details.
Figure 17: Visualization results of our DreamDrone. We generated 120 image sequences for each text prompt and visualized one image from every third frame to demonstrate the model’s capability in producing diverse and stable visual outputs over time.
Figure 18: Visualization results of our DreamDrone. We generated 120 image sequences for each text prompt and visualized one image from every third frame to demonstrate the model’s capability in producing diverse and stable visual outputs over time.
@@ 10 Limitation
11 Social impact
We introduce a new method for creating perpetual scenes from text descriptions, making it easier for people to generate high-quality images without needing complex training or data. This breakthrough can help in various areas, such as making educational content more engaging, aiding in environmental planning, and giving creative professionals new tools to express their ideas. As this technology becomes available, it’s important to use it wisely, ensuring it benefits society and does not contribute to misinformation or unethical use. In essence, DreamDrone offers exciting possibilities for innovation while emphasizing the need for responsible use.